40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 11 02:00:21 2013, Jenne, Update, Locking, TRY gone??? 
    Reply  Sun Apr 14 15:03:17 2013, Manasa, Update, Locking, Fixing - progress 
       Reply  Sun Apr 14 17:30:14 2013, Manasa, Update, Locking, Fixed 
          Reply  Sun Apr 14 23:20:42 2013, Den, Update, Locking, Fixed 
             Reply  Thu Apr 18 21:12:56 2013, Manasa, Update, Locking, Fixed 
Message ID: 8453     Entry time: Sun Apr 14 17:30:14 2013     In reply to: 8452     Reply to this: 8455  
Author: Manasa 
Type: Update 
Category: Locking 
Subject: Fixed 

TRY path fixed and ready for normalization.

I used 2" BS at R=50 and R=98 to reflect the Y arm transmission at QPD-Y and TRY PD respectively. The residual beam transmitted by the BS is now steered by a Y1mirror to the camera. With Y arm locked, transmission currently measures 40mW against the expected 70mW. TRY shows 0.45 counts in dataviewer.

ELOG V3.1.3-